OpenStudy (anonymous):

Math help!!!! pls Use 1% or 10% to estimate 59% of 80 Use a fraction to estimate 20% of 80 Rons parents deposited $3,000 into a savings account as a college fund when he was born. How much will Ron have in the account after 18 years at a yearly simple interest rate of 3.75%?

OpenStudy (anonymous):

can u find 20% of 80

OpenStudy (anonymous):

16

OpenStudy (anonymous):

6*10% of 80-1%of 80=6*8-0.8=47.2